tan y =√2 , where x and y are acute angles, find the value of x if sin x = 1 - cos y
1
Expert's answer
2019-09-11T10:36:01-0400

sin2y+cos2y=1; 1+tan2y=1/(cos2y); cos2y=1/(1+tan2y); cos y= (1/(1+tan2y))1/2; cos y = 3-1/2

sin x = 1 - 1/31/2= (31/2-1)/31/2; x = arcsin (31/2-1)/31/2

Using calculator x = 25 degrees